# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v1

@@ 3.1 The Challenge of Dynamic Re-Classification
To achieve this, we maintain a set of active large jobs, denoted AA, containing the $\ell$ jobs with the largest processing times seen so far. The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
Upon arrival, a job is evaluated. If it is not large enough to enter the top-$\ell$ group, it is permanently classified as a small job. Otherwise, it is designated as an active large job and added to the set AA.
@@
The main complexity arises when a job $j\in A$ is forced out (or retired) at time $r_{i}$ due to the arrival of a new, larger job ii. The state of $j$ at this moment determines its new classification:
@@
  • If job $j$ has already started processing by the online algorithm, it cannot be proxied, as the online algorithm must not revoke historical scheduling decisions. We call this a committed large job. Although no longer in the active set AA, it must still be tracked and scheduled as a large job. This is a subtle point in the analysis and can be safely ignored for a high-level understanding, but it is an important case required for the correctness of our online algorithm.
Data: Active large job set AA, and every job’s other current states
@@
We now present the pseudocode for this partitioning step in Algorithm 1. For simplicity, the function only explicitly manages the active large job set AA; the other sets (small, proxied, and committed jobs) are implicitly defined by the procedure. This partitioning mechanism guarantees our desired invariants along with a key monotonicity property: the set of all small jobs (including proxy jobs) forms an “easier” scheduling instance than the original job set. This is formalized in the following lemma.
@@
    The number of active large jobs (the set AA) is at most $\ell$ at any time.
@@ 3.2 Handling Small Jobs with Non-Preemptive Shortest Job First
We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every pp.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank ii.
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.
@@ 5 Deterministic Non-Preemptive Algorithm
  • Mixed machines (Machines 22 to $\lceil m/2\rceil+1$): Accept both small and large jobs.
@@ 6 Deterministic Algorithm with Kill-and-Restart
Data: Active large job set AA, and every job’s other current states
@@
  • Small-Only machines (Machines 11 to $\lfloor m/2\rfloor$): These machines accept only small jobs.
@@ 7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$. The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines 11 to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
@@
As a result, a job may switch between being large and small multiple times. For instance, when $n(t)$ is small, we may treat the top 33 jobs as large, so a job of rank 44 is small. Later, as more jobs arrive and $\ell(t)$ becomes 44, this job becomes large. If even more jobs arrive, $\ell(t)$ may increase to 55, but that same job may drop out of the top-55 and become small again. This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
@@
We use a partition rule that depends on the number of jobs released so far. Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use NN to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$. Let $T_{j}(t)$ denote the type of job $j$ at time $t$. We classify $T_{j}(t)$ as large if job $j$ is among the top $\ell$ jobs in size and $p_{j}\geq\frac{4P(t)}{\ell}$, where $P(t)$ is the total size of all jobs released by time $t$. Otherwise, $T_{j}(t)$ is classified as small.
A New Kill Event When NN Is Updated.
Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine. However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier. To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever NN is updated (that is, when $n(t)>N$, and we double NN to $2N$). At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so. We remark that this is one reason for using a rounded value of $n(t)$ (the nearest power of two) as the benchmark, since it allows us to bound the number of such kill events.
@@
 // Update the Partition of all jobs again based on the current number of jobs NN and the current sum of processing time $P$.
@@
 if NN is updated by the release of job $j$ then
@@
Then we bound $F_{j}^{(1)}$ and $F_{j}^{(2)}$ separately. The analysis for $F_{j}^{(1)}$ is similar to the case where $n$ is known. Note that we have two types of kill events. One occurs when NN is updated, which we call a type-A event; the other occurs when $\phi$ becomes large, which we call a type-B event. The number of type-B kill events can be bounded similarly to the known-$n$ case, with the help of the rounded benchmark value NN.
@@
We know that if a type-B kill event occurs at time $t$, there must be at least $\lfloor N(t)^{\alpha}\sqrt{m}\rfloor$ small jobs waiting, where $N(t)$ denotes the value of NN at time $t$. We interpret these small jobs as the creators of this kill event.
Note that when a small job $j$ creates a type-B kill event at time $t$, it remains small until NN is updated. Hence, after the kill event at $t$ and before the next update of NN, no large job will be processed on any machine unless we start job $j$. On the other hand, if we start $j$, it will not be killed again until NN is updated, since $j$ stays small during this period. Therefore, job $j$ cannot trigger another type-B kill event before NN changes.
Consequently, the number of kill events that occur while $N=x$ is at most $x^{1-\alpha}/\sqrt{m}$. Since NN increases geometrically by a factor of 22, the total number of kill events is bounded by $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right),$ as $N\leq 2n$. ∎
@@
Although NN may be smaller than $n$ since $n$ is unknown at that time, it is lower bounded by $n^{\alpha}\sqrt{m}/2$, because at least $n^{\alpha}\sqrt{m}$ small blocked jobs are waiting, implying $n(t)\geq n^{\alpha}\sqrt{m}$. Hence, the size of $j$ satisfies
@@
Case 2. Some Small-Only machines are not busy at the end. This only happens at a type-A kill event. If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment. Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when NN was updated to $2N$ — the only time a small job can become large. At such an update, at most NN jobs have been released, so there are at most NN blocked jobs. Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$. Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$ at most
@@
As NN doubles through $1,2,4,\ldots$ up to $n<2^{k}\leq 2n$, the total contribution sums to
@@ 8.1 Warm-Up: Randomized Lower Bound on a Single Machine
We start with the lower bound that any randomized non-preemptive online algorithm for a single machine is $\Omega(\sqrt{n})$ as a warm-up. The counterexample is constructed as Algorithm 6, which is a distribution of two possibilities: the only difference is whether the first $k$ $\varepsilon$-jobs are released at time 11 or 22.
@@
Release job 11 at time 0 with processing time 22 ;
@@
We analyze the counterexample described in Algorithm 6. For each of the two possibilities, the optimal algorithm can arrange job 11 at time 0 or 11 correctly and delay each $\varepsilon$-job by at most $O(k^{2}\varepsilon)$. Therefore, the optimal flow time is always $O(1)$. However, for any deterministic online algorithm, it has to decide whether to start job 11 at the beginning or not:
  • If the algorithm starts job 11 at time 0, then with probability $1/2$, the algorithm will delay the first $k$ $\varepsilon$-jobs by 11 for the case the randomized $\varepsilon$-jobs are released at time 11;
  • Otherwise, the algorithm cannot start job 11 before time $k+1$ for the case the randomized $\varepsilon$-jobs are released at time 22, or it will delay at least $k$ $\varepsilon$-jobs by 11.
@@ 8.2 Generalization to Multiple Machines
  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 1+(j(k+2))/k1+(j-(k+2))/{k} $1/k$
  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 2+(j(k+2))/k2+(j-(k+2))/{k} $1/k$
@@
  • A Large Job: Job 0, with processing time 22 and release time $t$.
@@
In each busy gadget, we can schedule the large job either at $5b+0$ or $5b+2$ depending on $c$, and every small job upon arrival. As a result, the total flow time in this gadget is at most 44 (the large job) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore $\textup{{OPT}}\leq 6mk$ by summing up all gadgets. ∎
@@
In a given batch, the set of mkmk jobs released in a time interval $[t,t+1)$ constitutes a small job period. A large job conflicts with this period if its execution starts within the interval [t1.5,t+0.5][t-1.5,t+0.5]. This definition guarantees that the large job’s execution (of length 2) overlaps with the period $[t,t+1)$ for a duration of at least $0.5$.
@@
Define $f(u)=m\cdot\lceil ku\rceil$ as the number of small jobs released at time $t+u$, and let $g(u)$ be the number of machines not processing large jobs at time $u$. The number of small jobs that can be completed by time $u$ is at most $\lfloor k\int_{0}^{u}g(y)dy\rfloor$. The total flow time of the kmkm small jobs in the batch is no less than
@@
If no large jobs conflict with this period, $g(u)$ is defined as $m$ for every $u\in[0,1)$. By the condition of the lemma, we have $x$ jobs that decrease $g(u)$ by 11 with a duration of at least $0.5$. The minimum contribution of the lower bound occurs when each large job conflicts the period at $[0.5,1)$, yielding
@@
Note that before time 11, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 11 is not related to the result of $c_{b}$.
  • Let $x$ be the number of large jobs started in [0.5, 1)[-0.5,\;1). If small jobs are Left, all these $x$ jobs contribute to II.
After time 11, the algorithm’s scheduling decisions may depend on $c_{b}$. We consider the two cases separately:

### v2

@@ 3.1 The Challenge of Dynamic Re-Classification
To achieve this, we maintain a set of active large jobs, denoted $A$, containing the $\ell$ jobs with the largest processing times seen so far. The logic of our framework defines a clear life cycle for every job based on its interaction with this set.
@@
Upon arrival, a job is evaluated. If it is not large enough to enter the top-$\ell$ group, it is permanently classified as a small job. Otherwise, it is designated as an active large job and added to the set $A$.
@@
The main complexity arises when a job $j\in A$ is forced out (or retired) at time $r_{i}$ due to the arrival of a new, larger job $i$. The state of $j$ at this moment determines its new classification:
@@
  • If job $j$ has already started processing by the online algorithm, it cannot be proxied, as the online algorithm must not revoke historical scheduling decisions. We call this a committed large job. Although no longer in the active set $A$, it must still be tracked and scheduled as a large job. This is a subtle point in the analysis and can be safely ignored for a high-level understanding, but it is an important case required for the correctness of our online algorithm.
Data: Active large job set $A$, and every job’s other current states
@@
We now present the pseudocode for this partitioning step in Algorithm 1. For simplicity, the function only explicitly manages the active large job set $A$; the other sets (small, proxied, and committed jobs) are implicitly defined by the procedure. This partitioning mechanism guarantees our desired invariants along with a key monotonicity property: the set of all small jobs (including proxy jobs) forms an “easier” scheduling instance than the original job set. This is formalized in the following lemma.
@@
    The number of active large jobs (the set $A$) is at most $\ell$ at any time.
@@ 3.2 Handling Small Jobs with Non-Preemptive Shortest Job First
We discuss two timepoints $t+\tau$ and $t+2\tau$. By Lemma 3, we have the volume bound: $V^{\prime}_{\leq p}(t+\tau)\geq V_{\leq p}(t)$ for every $p$.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t+\tau$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all rank $i$.
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t+\tau)<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 3. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
@@ 5 Deterministic Non-Preemptive Algorithm
  • Mixed machines (Machines $2$ to $\lceil m/2\rceil+1$): Accept both small and large jobs.
@@ 6 Deterministic Algorithm with Kill-and-Restart
Data: Active large job set $A$, and every job’s other current states
@@
  • Small-Only machines (Machines $1$ to $\lfloor m/2\rfloor$): These machines accept only small jobs.
@@ 7 Deterministic Algorithm with Kill-and-Restart without the Knowledge of $n$
In this section, we extend the $\sqrt{n/m}$-competitive algorithm (which assumes the prior knowledge of $n$) to an $O(n^{\alpha}/\sqrt{m})$-competitive algorithm without such knowledge, where $\alpha=(\sqrt{5}-1)/2$. The overall framework still follows Section 6: we keep the same machine partition (Small-Only machines $1$ to $\lfloor m/2\rfloor$ and Mixed machines $\lfloor m/2\rfloor+1$ to $m$), and we still kill a large job once enough small jobs are blocked.
@@
As a result, a job may switch between being large and small multiple times. For instance, when $n(t)$ is small, we may treat the top $3$ jobs as large, so a job of rank $4$ is small. Later, as more jobs arrive and $\ell(t)$ becomes $4$, this job becomes large. If even more jobs arrive, $\ell(t)$ may increase to $5$, but that same job may drop out of the top-$5$ and become small again. This uncontrolled oscillation between the large and small states makes the partition unstable and invalidates the previous analysis.
@@
We use a partition rule that depends on the number of jobs released so far. Instead of using $n(t)$ directly, we let $N=2^{k}$ be the smallest power of two such that $N\geq n(t)$, and use $N$ to define the large–small threshold $\ell$ at every time $t$, where $\ell=\lfloor N^{\alpha}\sqrt{m}\rfloor$. Let $T_{j}(t)$ denote the type of job $j$ at time $t$. We classify $T_{j}(t)$ as large if job $j$ is among the top $\ell$ jobs in size and $p_{j}\geq\frac{4P(t)}{\ell}$, where $P(t)$ is the total size of all jobs released by time $t$. Otherwise, $T_{j}(t)$ is classified as small.
A New Kill Event When $N$ Is Updated.
Previously, we triggered a kill event only when too many small jobs were waiting while a large job was being processed on a Mixed machine. However, in this setting, a large job may be processed on a Small-Only machine, since it might have been classified as small earlier. To prevent this from causing inefficiency, we introduce a new kill event that occurs whenever $N$ is updated (that is, when $n(t)>N$, and we double $N$ to $2N$). At this moment, we check whether any currently large jobs are running on the Small-Only machines and kill all of them if so. We remark that this is one reason for using a rounded value of $n(t)$ (the nearest power of two) as the benchmark, since it allows us to bound the number of such kill events.
@@
 // Update the Partition of all jobs again based on the current number of jobs $N$ and the current sum of processing time $P$.
@@
 if $N$ is updated by the release of job $j$ then
@@
Then we bound $F_{j}^{(1)}$ and $F_{j}^{(2)}$ separately. The analysis for $F_{j}^{(1)}$ is similar to the case where $n$ is known. Note that we have two types of kill events. One occurs when $N$ is updated, which we call a type-A event; the other occurs when $\phi$ becomes large, which we call a type-B event. The number of type-B kill events can be bounded similarly to the known-$n$ case, with the help of the rounded benchmark value $N$.
@@
We know that if a type-B kill event occurs at time $t$, there must be at least $\lfloor N(t)^{\alpha}\sqrt{m}\rfloor$ small jobs waiting, where $N(t)$ denotes the value of $N$ at time $t$. We interpret these small jobs as the creators of this kill event.
Note that when a small job $j$ creates a type-B kill event at time $t$, it remains small until $N$ is updated. Hence, after the kill event at $t$ and before the next update of $N$, no large job will be processed on any machine unless we start job $j$. On the other hand, if we start $j$, it will not be killed again until $N$ is updated, since $j$ stays small during this period. Therefore, job $j$ cannot trigger another type-B kill event before $N$ changes.
Consequently, the number of kill events that occur while $N=x$ is at most $x^{1-\alpha}/\sqrt{m}$. Since $N$ increases geometrically by a factor of $2$, the total number of kill events is bounded by $O\!\left(\frac{n^{1-\alpha}}{\sqrt{m}}\right),$ as $N\leq 2n$. ∎
@@
Although $N$ may be smaller than $n$ since $n$ is unknown at that time, it is lower bounded by $n^{\alpha}\sqrt{m}/2$, because at least $n^{\alpha}\sqrt{m}$ small blocked jobs are waiting, implying $n(t)\geq n^{\alpha}\sqrt{m}$. Hence, the size of $j$ satisfies
@@
Case 2. Some Small-Only machines are not busy at the end. This only happens at a type-A kill event. If a blocked job is waiting, it is in its last small period, so it is small and all machines are busy at that moment. Hence, if some Small-Only machines are idle at the end, a type-A kill event must have occurred when $N$ was updated to $2N$ — the only time a small job can become large. At such an update, at most $N$ jobs have been released, so there are at most $N$ blocked jobs. Each killed job was small just before the update, hence its size is at most $\textup{{OPT}}/(N^{\alpha}\sqrt{m})$. Thus each such event contributes to $\sum_{h\in H_{2}}\sum_{j\in B_{h}}F^{(2)}_{j}$ at most
@@
As $N$ doubles through $1,2,4,\ldots$ up to $n<2^{k}\leq 2n$, the total contribution sums to
@@ 8.1 Warm-Up: Randomized Lower Bound on a Single Machine
We start with the lower bound that any randomized non-preemptive online algorithm for a single machine is $\Omega(\sqrt{n})$ as a warm-up. The counterexample is constructed as Algorithm 6, which is a distribution of two possibilities: the only difference is whether the first $k$ $\varepsilon$-jobs are released at time $1$ or $2$.
@@
Release job $1$ at time 0 with processing time $2$ ;
@@
We analyze the counterexample described in Algorithm 6. For each of the two possibilities, the optimal algorithm can arrange job $1$ at time 0 or $1$ correctly and delay each $\varepsilon$-job by at most $O(k^{2}\varepsilon)$. Therefore, the optimal flow time is always $O(1)$. However, for any deterministic online algorithm, it has to decide whether to start job $1$ at the beginning or not:
  • If the algorithm starts job $1$ at time 0, then with probability $1/2$, the algorithm will delay the first $k$ $\varepsilon$-jobs by $1$ for the case the randomized $\varepsilon$-jobs are released at time $1$;
  • Otherwise, the algorithm cannot start job $1$ before time $k+1$ for the case the randomized $\varepsilon$-jobs are released at time $2$, or it will delay at least $k$ $\varepsilon$-jobs by $1$.
@@ 8.2 Generalization to Multiple Machines
  Left (when $c=0$) $\{k+2,\dots,2k+1\}$ 1+(i(k+2))/k1+(i-(k+2))/{k} $1/k$
  Right (when $c=1$) $\{k+2,\dots,2k+1\}$ 2+(i(k+2))/k2+(i-(k+2))/{k} $1/k$
@@
  • A Large Job: Job 0, with processing time $2$ and release time $t$.
@@
In each busy gadget, we can schedule the large job either at $5b+0$ or $5b+2$ depending on $c$, and every small job upon arrival. As a result, the total flow time in this gadget is at most $4$ (the large job) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore $\textup{{OPT}}\leq 6mk$ by summing up all gadgets. ∎
@@
In a given batch, the set of $mk$ jobs released in a time interval $[t,t+1)$ constitutes a small job period. A large job conflicts with this period if its execution starts within the interval [t1.5,t+0.5][t-1.5,t+0.5]. This definition guarantees that the large job’s execution (of length 2) overlaps with the period $[t,t+1)$ for a duration of at least $0.5$.
@@
Define $f(u)=m\cdot\lceil ku\rceil$ as the number of small jobs released at time $t+u$, and let $g(u)$ be the number of machines not processing large jobs at time $u$. The number of small jobs that can be completed by time $u$ is at most $\lfloor k\int_{0}^{u}g(y)dy\rfloor$. The total flow time of the $km$ small jobs in the batch is no less than
@@
If no large jobs conflict with this period, $g(u)$ is defined as $m$ for every $u\in[0,1)$. By the condition of the lemma, we have $x$ jobs that decrease $g(u)$ by $1$ with a duration of at least $0.5$. The minimum contribution of the lower bound occurs when each large job conflicts the period at $[0.5,1)$, yielding
@@
Note that before time $1$, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time $1$ is not related to the result of $c_{b}$.
  • Let $x$ be the number of large jobs started in [0.5, 1)[-0.5,\;1). If small jobs are Left, all these $x$ jobs contribute to $I$.
After time $1$, the algorithm’s scheduling decisions may depend on $c_{b}$. We consider the two cases separately: